OpenStudy (anonymous):

8. What is the solution to the equation √(2x+8 )-6=4

OpenStudy (dejavo):

x=15

OpenStudy (anonymous):

how u gott x=15 @dejavo

OpenStudy (haseeb96):

√(2x+8 )-6=4 √(2x+8 )=4+6 √(2x+8 )=10 taking square on both sides then 2x + 8=10 x 10 2x+8=100 2x=100-8 2x=92 x=92\2 x=46 verification put the value of x in given equation √(2(46)+8 )-6=4 \[\sqrt{92+8}-6=4\] \[\sqrt{100}-6=4\] 100 is square of 10 10-6=4 4=4 Hence proved x=46 This is the correct answer

OpenStudy (anonymous):

x=1